OpenStudy (anonymous):

x+4z=17; -3x+2y-z=-20; x-5y+3z=19

OpenStudy (anonymous):

alright so take your first equation solve it for x \[x=17-4z\] (all i did was subtract 4z on both sides) take this and put it into the 3rd equation for x \[(17-4z)-5y+3z=19\] \[17-4z-5y+3z=19\] \[17-z-5y=19\] solve this for z \[(17-19)-5y=z\] \[-2-5y=z\]

OpenStudy (anonymous):

lastly take what you got from x and z and put that into the 2nd equation \[-3(17-4(-2-5y))+2y-(-2-5y)=-20\]

OpenStudy (anonymous):

now simplify this and you'll get an equation that you can solve for y, which you can use to get z and use what you got for z to get x